OpenStudy (anonymous):

For f(h)=-6h-2/3, find f(-3)

OpenStudy (anonymous):

16\3

OpenStudy (anonymous):

no not that haha

OpenStudy (gorv):

f(-3)=-6*(-3)-2/3=18-2/3=(54-2)/3=52/3

OpenStudy (anonymous):

hrmmm.. thats not on my answer choice sorry :(

OpenStudy (anonymous):

a. 29 1/3 b. 17 1/3 c.-3 2/3 d.-24 2/3

OpenStudy (gorv):

17 1/3